Pre-Algebra: Find the three solutions to the equation below.
2x - 4y = 4
a. (3,0) (4,0 (-4,0)
b. (2,0) (8,4) (-2,1)
c. (0,-1) (2,0) (1,-1/2)

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

Let's go through the choices

Choice A) (3,0) (4,0) (-4,0)

the first point of choice A is (3,0) which means x = 3 and y = 0. They BOTH need to make the equation true.

2x - 4y = 4
2(3) - 4(0) = 4 ... plug in (x,y) = (3,0)
6 - 0 = 4
6 = 4 ... this is false, so (3,0) is NOT a solution making choice A NOT an answer. We move on to the next choice

----------------------------------------------------------------

Choice B) (2,0) (8,4) (-2,1)


Checking (2,0)

2x - 4y = 4
2(2) - 4(0) = 4
4 - 0 = 4
4 = 4 ... looks good

Checking (8,4)

2x - 4y = 4
2(8) - 4(4) = 4
16 - 16 = 4
0 = 4 ... false, so we eliminate choice B

----------------------------------------------------------------
Choice C) (0,-1) (2,0) (1,-1/2)

Checking (0,-1)

2x - 4y = 4
2(0) - 4(-1) = 4
0 + 4 = 4
4 = 4 ... looks good

Checking (2,0)

2x - 4y = 4
2(2) - 4(0) = 4
4 - 0 = 4
4 = 4 ... looks good

Checking (1, -1/2)

2x - 4y = 4
2(1) - 4(-1/2) = 4
2 + 2 = 4
4 = 4 ... looks good

==========================================================================

So in the end, choice C is the answer since ALL of the points listed in choice C make the equation true.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------

Answer by stanbon(75887)

Find the three solutions to the equation below.
2x - 4y = 4
---
Solve for "Y":
y = (1/2)x - 1
------
a. (3,0)no (4,0)no (-4,0)no
b. (2,0)yes (8,4)no (-2,1)no
c. (0,-1)yes (2,0)yes (1,-1/2)yes
--------
Cheers,
Stan H.
===========